# Paper/main_online_appendix

### v3

@@ 4.2 Tests of Preference Reversal Predictions
We run experiments mirroring the simulation exercises in Section 3.3 to show that preference reversals can be explained as an outgrowth of comparison complexity. Specifically, we establish the presence of reversals in lottery and intertemporal choice, and test the novel model prediction that these reversals can be eliminated by manipulating the ease of comparing options to the price list.
In each domain, we construct two sets of “base" options: 3 high-risk and 3 low-risk lotteries, and 3 high-delay and 3 low-delay payoff flows (see Table 2). We elicit binary choices between these sets of options, as well as multiple price list valuations of each option. In these valuation tasks, we experimentally vary the ease of comparing each option to the price list: for lotteries, we elicit certainty equivalents as well as probability equivalents using a yardstick lottery that pays \$24 with $p$chance; for intertemporal choice, we elicit present value equivalents as well as time equivalents using a yardstick option that pays \$27.5 in $t$ days. We elicit choices and valuations for both the base options as well as a “scaled-up” version of each option, where we multiply payouts by a scale factor of 1.6.
@@
Figures 8 and 9 present results from our lottery and intertemporal choice experiments, aggregating across scale factors. Figures 8(a) and 9(a) show that binary choice rates favor the low-risk and low-delay options, with choice rates for those options well above 50%. In contrast, Figures 8(b) and 9(b) show that there is an apparent reversal in preference when the options are valued via certainty equivalents and present value equivalents: subjects assign higher valuations to high-risk and high-delay options on average. In these same figures, however, we see the predicted flipping of these valuation patterns when we manipulate the ease of comparing each option to the price list. In particular, when options are instead valued via probability equivalents and time equivalents, subjects instead assign higher valuations to the low-risk and low-delay options, thus eliminating the apparent reversal. [^25] Butler and Loomes, (2007) document a related result in lottery choice: examining inconsistency between direct choice and probability equivalents, they find that subjects are more likely to state higher PEs for the low-risk lottery yet choose the high-risk lottery in direct choice than to exhibit the opposite inconsistency. We find similar patterns in our data.
@@ 4.3 Tests of Valuation Predictions
We run a multiple price list experiment to 1) replicate familiar patterns of inverse-S shaped probability weighting and hyperbolic discounting in certainty equivalents and present value equivalents respectively, and 2) test the prediction, developed in Section 3.4, that these biases can be reversed by manipulating the units of valuation. Specifically, our model predicts that relative to the certainty equivalents of binary lotteries, probability equivalents of certain payments should exhibit underweighting of low probabilities and overweighting of high probabilities. Likewise, relative to the discounting revealed by the present value equivalents of delayed payments, time equivalents of immediate payments should exhibit overvaluation of delays close to the present and undervaluation of delays far in the future.
@@
Lottery Valuations. We elicit certainty equivalents $CE(l)$ for a simple lottery $l=(\overline{w},p_{l})$, and relate the normalized valuations $CE(l)/\overline{w}$ to payout probabilities $p_{l}$. We also elicit probability equivalents $PE(c)$ of a certain payment $c=(w_{c},1)$ against a yardstick lottery $(\overline{w},p)$, and relate the normalized payouts $w_{c}/\overline{w}$ to probability equivalents $PE(c)$. In our experiment, we draw $\overline{w}$ from {\$$\{\$9,\$18,\$27\}$. For certainty equivalents, we draw $p_{l}$ from $\{0.03,0.05,0.10,0.25,0.5,0.75,0.90,0.95,0.97\}$. For probability equivalents, we draw $w_{c}$ so that $w_{c}/\overline{w}\in\{0.033,0.056,0.11,0.25,0.5,0.75,0.89,0.944,0.967\}$.
@@
Results for lotteries are presented in Figure 10. The probability weighting function implied by certainty equivalents follows the familiar pattern of overweighting small probabilities and underweighting large probabilities. Consistent with the model’s predictions, the pattern reverses for probability equivalents. Here, small probabilities are instead underweighted, and large probabilities are overweighted. [^26] These results are consistent with Bouchouicha et al., (2023), who find similar patterns in probability equivalents, although their data does not extend to the range of normalized certain payments for which we find that subjects appear risk-seeking $(w_{c}/\overline{w}\geq 0.89)$.
@@ 4.4 Benchmarking Model Performance
Whereas the predominant approach in behavioral economics is to model non-standard elements in the DM’s value function, this paper seeks to model an orthogonal feature of choice: how difficult the DM finds a choice, given her objectives. To study whether our theory indeed explains variation in choice patterns that are not captured by existing models, we compare the explanatory power of these modeling approaches in our binary choice data.
First, we structurally estimate our choice model, which is parameterized by a transformation $G$ that maps the value-dissimilarity ratio into choice probabilities and (in the case of lottery and intertemporal choice) domain-specific preference parameters. Across all domains, we use the same specification of $G$ developed in Section 2.4. We then compare performance against a set of models specifying the DM’s value function, which we fit to our data using logit errors. In particular, we estimate a “standard" benchmark model: a distortion-free logit model in multiattribute choice, exponential discounted utility in intertemporal choice, and expected utility in lottery choice. We also estimate a set of behavioral models: salience-weighting (Bordalo et al.,, 2013), focusing (Kőszegi and Szeidl,, 2013), and relative thinking (Bushong et al.,, 2021) in multiattribute choice; quasi-hyperbolic and hyperbolic discounting in intertemporal choice; and cumulative prospect theory in lottery choice. Estimates are obtained using maximum likelihood; see Appendix E for details on each specification.
@@
Figure 12: Variance explained of estimated models and completeness benchmarks. $R^{2}$ values are observation-weighted. Number of free model parameters in parentheses. “LGT”, “RT”, “L1-C” refer to the Distortion-Free, Relative Thinking, and $L_{1}$-complexity models described in Appendix E.1. “EDU”, “HDU”, and “CPF-C” refer to the Exponential Discounting, Hyperbolic Discounting, and CPF-complexity models desdribed in Appendix E.2. ‘EU", “CPT", “EV CDF-C", and “EU CDF-C" refer to the Expected Utility, Cumulative Prospect Theory, risk-neutral CDF-complexity, and expected utility CDF-complexity models described in Appendix E.3. Completeness benchmarks are obtained using an ensemble of fitted models and a neural network; see Online Appendix G for details.
@@
Completeness and Restrictiveness. Following Fudenberg et al., (2022), we establish “completeness benchmarks" in each of our three domains by training flexible predictive models to predict choice rates based on problem features; these benchmarks represent the predictable variation in choice rates that any model could hope to capture. We use an ensemble approach to form these benchmarks, described in detail in Online Appendix G, where we combine parametric model predictions with those of a neural network trained on the data. The dashed lines in Figure 12 report the $R^{2}$ of these benchmarks, and Appendix Tables 5, 8, and 11 report the completeness index of each model following Fudenberg et al., (2022). Our model captures 70% of the predictable variation in our multiattribute choice data, and over 90% in lottery and intertemporal choice.
@@
One concern is that the predictive power of our model may come at the cost of parsimony — i.e. our model is flexible enough to explain a wide range of counterfactual datasets. To address this, we estimate the restrictiveness of each model: a measure of how well it fits synthetic data. Following the methodology developed in Fudenberg et al., (2023), we simulate synthetic data in each domain according to a set of common restrictions imposed by the models we evaluate, and assess how well each model fits the synthetic data on average (see Appendix F for details). We obtain a restrictiveness measure $r\in[0,1]$, where $r=1$ means the model fits the synthetic data no better than a fixed baseline model on average, and $r=0$ means the model perfectly fits the synthetic data. Appendix Tables 5, 8, and 11 report this measure in each domain. While our model is less restrictive than relative thinking in multiattribute choice ($r=0.45$ vs. 0.46), our model is more restrictive than cumulative prospect theory ($r=0.60$ vs. 0.53) and hyperbolic discounting ($r=0.59$ vs. 0.58) in lottery and intertemporal choice, respectively. These results suggest that the performance gains of our model do not come at substantial costs to model parsimony.
5 Strategic Obfuscation
We now apply our model to study strategic obfuscation in markets. Our application is motivated by the practice of product proliferation, in which firms market an otherwise undifferentiated good in a large number of different packaging varieties in which quantities and other peripheral characteristics differ.
(a) Advil Ibuprofen tablet varieties
(b) Bounty paper towel varieties
Figure 13: Examples of product proliferation.
To illustrate, consider Figure 13(a), which shows a subset of Amazon search results for Advil tablets. Notice that across these listings it is not simply the total quantity that varies — quantities are expressed multidimensionally. For instance, the first two listings offer the same total quantity, marketed as 1 package of 100 tablets one case and 50 packages of 2 tablets in the other. The third and fourth listings market 2 packages of 24 tablets vs. 1 package of 50 tablets, again for a similar total quantity. Analogous multidimensional product proliferation can be found in a number of markets, such as for paper towels, where the size and number of rolls vary (see Figure 13(b)), or for soft drinks, where the size and number of cans in a package vary.
A standard explanation for this practice is that firms are catering to consumers with heterogenous preferences over how quantities are packaged. However, an alternative explanation has been discussed in the literature: product proliferation may be a strategic choice to make comparison shopping difficult, thus reducing consumers’ price sensitivity (Ellison and Ellison,, 2018; Richards et al.,, 2020). Writing on the Ibuprofen market, Ellison and Ellison, (2018) remark “Sorting on price would be meaningless because these products, all with a chemically identical active ingredient, differ on so many dimensions, such as strength of tablet, size of package, and type of packaging.” In the following section, we apply our model to formally study how comparison complexity can generate this kind of spurious differentiation in equilibrium; in Section 5.2, we return to the question of how the analyst may distinguish between consumer heterogeneity and confusion.
5.1 Duopoly with Strategic Obfuscation
5.1.1 Symmetric Duopoly
We consider a two-stage game in which identical firms first choose a multidimensional product quantity vector, and then play a simultaneous Bertrand pricing game. We begin by analyzing the second stage game in which firms take quantities as given.
Pricing Game. Each firm $i\in\{a,b\}$ sells a good with multidimensional quantity $q^{i}=(q^{i}_{1},q^{i}_{2})$. We abstract away from vertical quantity differentiation and assume that firms sell the same total quantity, i.e. $q_{1}^{i}+q_{2}^{i}=Q$ for $i=a,b$. Firms face identical marginal costs $c$ for producing a unit of quantity $Q$. Each firm takes quantities as given and simultaneously chooses a price $p^{k}$; let $i=(p^{i},q^{i}_{1},q^{i}_{2})$ denote firm $i$’s offering.
There is a unit mass of identical consumers with unit demand, with utility given by $U(i)=q^{i}_{1}+q^{i}_{2}-p^{i}$. Consumers behave according to the $L_{1}$-complexity model, wherein choice probabilities are given by
$\rho(i,-i)=G\left(\frac{U(i)-U(-i)}{d_{L1}(i,-i)}\right)$
for $G$ strictly increasing, where we additionally assume that $G(1)=1,G(-1)=0$, so that consumers do not err in the face of dominance. Each firm’s profits are given by
$\Pi(i,-i)=(p^{i}-c)G\left(\frac{U(i)-U(-i)}{d_{L1}(i,-i)}\right)$
We analyze the pure-strategy Nash equilibria of the simultaneous pricing game.
Proposition 3.
If $G$ is linear, then there is a unique pure-strategy equilibrium, in which $p^{a}=p^{b}=c+\Delta q$, where $\Delta q=|q^{a}_{1}-q^{b}_{1}|+|q^{a}_{2}-q^{b}_{2}|$ denotes the dissimilarity between the firms’ quantity profiles. More generally, if a symmetric pure-strategy equilibrium exists, it is unique and satisfies $p^{a}=p^{b}=c+\frac{1}{2G^{\prime}(0)}\Delta q$.
Note that in the frictionless Bertrand benchmark in which consumers choose the offering that maximizes their utility, firms price at marginal cost and earn zero profit in equilibrium. Proposition 3 states that due to complexity-driven errors, markups can be sustained in equilibrium so long as firms offer dissimilar quantity profiles, where the magnitude of the markups is increasing in the degree of dissimilarity. [^30] Of course, a number of well-studied oligopoly models predict equilibrium markups with frictionless consumers (e.g. Kreps and Scheinkman,, 1983). We focus here on a Bertrand game to isolate the effect of comparison complexity from these other mechanisms. Richards et al., (2020) provides suggestive evidence for this prediction in the soft drink market: focusing on the different packaging varieties sold by retailers (chiefly characterized by the number of cans and the size of each can), they find that within a soft drink brand, package varieties that are unique to a single retailer are sold at higher markups. [^31] Using a structural model, the authors argue that a number of standard explanations, such as preference heterogeneity, cannot account for these patterns, and interpret them as strategic obfuscation by retailers.
This paper is not the first to formalize the idea that complexity can lead to higher markups by softening price competition. A line of theoretical work has studied how complexity-driven decision noise produces equilibrium markups, and how firms in turn may strategically induce complexity (Carlin,, 2009; Piccione and Spiegler,, 2012; Chioveanu and Zhou,, 2013; Gabaix et al.,, 2016). This literature has proceeded by studying market complexity in a reduced-form way: for instance, Carlin, (2009) model complexity as a scalar chosen by firms to influence the share of noisy consumers. Our primary contribution is to microfound one source of market complexity — the difficulty of trading off product attributes — and thereby relate choice complexity to observable product characteristics. [^32] One paper that takes a similar approach is Bachi and Spiegler, (2018), which analyzes a duopoly model in which consumers resort to a fixed random choice rule when faced with options that do not have a dominance relationship. In their model, all prices and product characteristics are chosen simultaneously, leading to a different set of predictions. This produces predictions about the equilibrium patterns of product characteristics, as we discuss below.
Endogenous Quantities. We now consider the first-stage game, in which firms choose their quantity profiles. In particular, consider a setting where each firm simultaneously [^33] The equilibria in Corollary 3.1 are unchanged if instead firms move sequentially in the first-stage game. chooses a quantity profile along a segment of an indifference curve in quantity space: $(q_{1},q_{2})\in\{[\underline{q},\overline{q}]^{2}:q_{1}+q_{2}=Q\}$, where $\overline{q}+\underline{q}=Q$.
Assume that $G$ is linear. Since in the subsequent pricing game, each firm’s profits are given by $\frac{1}{2}\Delta q$, firms will want to locate as far apart from each other as possible.
Corollary 3.1.
Suppose $G$ is linear. A pure-strategy equilibrium of the first-stage game exists, and in any such equilibrium, the equilibrium quantity dissimilarity is $\Delta q=2(\overline{q}-\underline{q})$.
This echoes the basic result from the classic Hotelling model: firms are driven to differentiate from their competitors to soften price competition. The key difference lies in the welfare interpretation of this differentiation: in the Hotelling model, differentiation caters to real taste differences between consumers, and so the efficient level of differentiation is in general positive. In our setting, however, differentiation is entirely spurious, and results purely in a transfer from firms to consumers.
5.1.2 Imitation vs. Obfuscation
As shown above, in a symmetric environment firms are driven to design products that are difficult to compare to their competitors. Below, we show that allowing for vertical cost differences between firms leads to a richer set of equilibrium predictions, in which firms may imitate their competition to increase price sensitivity.
Consider a setting identical to the one studied above, except that firms have different costs. In particular, let $c^{a}$ and $c^{b}$ denote the marginal costs for the two firms, and assume that $a$ is the lower cost firm; Let $\Delta c=c^{b}-c^{a}$ denote $a$’s cost advantage.
Rather than a simultaneous-move first-stage game, we consider a sequential first-stage game in which firm $b$ first announces a location decision in $\{(q_{1},q_{2})\in[\underline{q},\overline{q}]^{2}:q_{1}+q_{2}=Q\}$, followed by firm $a$ [^34] In a simultaneous first-stage game, pure-strategy equilibria will not exist if $a$ wishes to imitate, following the logic of hide-and-seek games. As before, profits are determined by a simulataneous-move pricing game played after the first-stage game. To ensure that pure-strategy equilibria are defined, we assume that if $q^{a}=q^{b}$ and $p^{a}=p^{b}$, $\rho(a,b)=\lim_{p\uparrow p^{a}}\rho((p,q^{a}),b)$ – that is, ties are broken in favor of the lower-cost firm if both firms sell identical products at identical prices.
As before let $\Delta q=|q_{1}^{a}-q_{1}^{b}|+|q_{2}^{a}-q_{2}^{b}|$ denote the quantity dissimilarity between the firms’ offerings, and let $\overline{\Delta q}=2(\overline{q}-\underline{q})$ denote the maximal quantity dissimilarity that can be implemented.
Proposition 4.
Suppose $G$ is linear, and restrict to pure-strategy equilibria.
  1. 1.
    If $\Delta c>2\overline{\Delta q}$, firm $a$ imitates, i.e. $\Delta q=0$ in equilibrium. The unique equilibrium prices satisfy $p^{a}=p^{b}=c^{b}$, and $b$’s equilibrium market share is 0.
  2. 2.
    If $\Delta c<2\overline{\Delta q}$, firm $a$ obfuscates, i.e. $\Delta q=\overline{\Delta q}$ in equilibrium. The unique equilibrium prices satisfy $p^{a}=c^{b}+\overline{\Delta q}$, $p^{b}>p^{a}$. Furthermore, $b$’s equilibrium market share is positive and strictly increasing in $\overline{\Delta q}$.
The intuition is as follows. Given the high-cost firm’s product choice, the low cost firm can pursue two strategies: they can either imitate their competitor, designing an identical product and using their cost advantage to capture the entire market, or they can obfuscate, designing a product that is difficult to compare and pricing high. [^35] If instead the low-cost firm $a$ moves first in the first-stage game, the minimum dissimilarity $a$ can implement is $1/2\overline{\Delta q}$. In this case, equilibria will either involve full obfuscation ($\Delta q=\overline{\Delta q}$) as before, or partial obfuscation $(\Delta q=1/2\overline{\Delta q})$, with some threshold $c^{*}$ such we have full obfuscation for $c<c^{*}$ and partial obfuscation otherwise. The relative attractiveness of these strategies depends on how large the firm’s cost advantage is relative to the extent to which obfuscation is possible. Here, obfuscation generates market inefficiency: whereas imitation leads to the efficient allocation in which the low-cost firm serves the entire market, obfuscation allows the high-cost firm to sustain positive market share.
Note one key comparative static prediction: as $a$’s cost advantage decreases, obfuscation becomes a more attractive strategy. As such, the model predicts that an increase in competitive pressure brought on by a decrease in $b$’s costs can actually lead to an increase in equilibrium prices and obfuscation.
5.2 Confusion vs. Heterogeneity
A standard explanation for product proliferation is that it simply reflects heterogeneous tastes over packaging varieties. Indeed, the binary choice probabilities in the $L_{1}$-complexity model are observationally equivalent to the choice shares generated by a population of agents endowed with a distribution over linear attribute weights $\beta$, where each agent chooses correctly. [^36] Formally, the $L_{1}$-complexity model shares an intersection with pure characteristics models (Berry and Pakes,, 2007); see Appendix B.8. These two interpretations of the data have dramatically different welfare and policy implications — for instance, policies that standardize product attributes may make sense if low price sensitivity reflects confusion, but not if it reflects heterogeneous tastes. As such, it is important to understand the relative importance of each in any given application. Here, we outline two ways in which these accounts may be distinguished.
First, individual-level choice data can be used to distinguish between heterogeneity and confusion. If population-level choice data reflects only taste heterogeneity, we should not expect within-individual choices to be noisy or inconsistent. In particular, Theorem 1 implies that with sufficient within-individual choice data, the analyst can identify both the individual’s utility (parameterized by $\beta$), as well as the extent to which the individual choices are prone to complexity-driven errors (parameterized by $G$). In principle, this allows the analyst to jointly estimate the degree of true preference heterogeneity and confusion within a population. As likely few observational datasets are sufficiently rich for such an exercise, this points to the potential importance of survey data. In particular, one might adapt conjoint surveys, which are traditionally used by researchers and practitioners to study consumer preferences, to also shed light on the degree of complexity-driven confusion.
Second, the observational equivalence between $L_{1}$-complexity and taste heterogeneity holds only for binary choice data. As discussed in Section 3.2.1, our multinomial choice model predicts violations of regularity, wherein the addition of choice options can reduce the choice share of existing options. This feature can not only be used to distinguish between complexity and preference heterogeneity, but itself generates novel predictions on market outcomes, as the following example illustrates.
Example 4.
(Three Firms.) Consider the duopoly pricing game analyzed in Section 5.1.1: we have firms $a,b$ with marginal costs $c$ selling the same total quantity. As before, let $\Delta q$ denote the quantity dissimilarity between the two firms. Consider the addition of a third firm $s$, which has identical quantity attributes as firm $a$: $(q^{a}_{1},q^{a}_{2})=(q^{s}_{1},q^{s}_{2})$, but with costs $c^{s}>c$. Binary choice probabilities are as described in the $L_{1}$-complexity model: $\rho(i,j)=G\left(\frac{U(i)-U(j)}{d_{L1}(i,j)}\right)$, where we also assume that $G$ is linear, i.e. $G(r)=\frac{1}{2}+\frac{1}{2}r$. We analyze the pure-strategy equilibria of the simultaneous-move pricing game.
First, consider the case where $c^{s}>c+\Delta q$: that is, firm $s$ would be priced out by $a$ under the equilibrium duopoly prices $p^{a}=p^{b}=c+\Delta q$ (see Proposition 3). Note that if binary choice probabilities reflect preference heterogeneity, the addition of $s$ would have no impact on the equilibrium market outcome: $a$ and $b$ maintain their duopoly prices, and $s$ is priced out of the market and so has no impact on demand. Under our multinomial choice model, however, the addition of $s$ will have an impact on demand, even if it is priced out by $a$: in particular, the fact that $s$ is dominated by $a$ but not by $b$ increases $a$’s market share (see Proposition 1), which then affects equilibrium prices.
To facilitate closed-form analysis of this equilibrium, we consider a binary-signal version of our multinomial choice model: for each firm $i,j$, consumers receive a signal that reveals the ordinal ranking between $i,j$ with probability
$\tau_{ij}=H\left(\frac{|U(i)-U(j)|}{d_{L1}(i,j)}\right)$
and contains no information otherwise. [^37] It can be shown that the binary-signal version of the model is observationally equivalent to the Gaussian-signal model in binary choice, and that all formal results and qualitative simulation results in Section 3 hold in this version of the model. Let $\rho(p^{i},\{p^{i},p^{j},p^{k}\})$ denote the demand for $i$’s offering as a function of the prices of the three firms. To ensure that pure strategy equilibria exist, we assume the same tiebreaking rule as in Section 5.1.2: in the case where firms $a$ and $s$ provide an identical offering, we break ties in favor of $a$: $\rho(p^{a},\{p^{a},p^{b},p^{s}\})=\lim_{p\uparrow p^{a}}\rho(p,\{p,p^{b},p^{c}\})$ whenever $p^{a}=p^{s}$.
Proposition 5.
Suppose $H$ is linear, i.e. $H(r)=r$.
  1. 1.
    A pure-strategy equilibrium exists, and in any such equilibrium, $p^{b}<p^{a}$.
  2. 2.
    If $\Delta q\geq c_{s}-c$, there is a unique pure-strategy equilibrium in which $s$ prices at cost, and $b$’s equilibrium profit is decreasing in $\Delta q$.
The intuition is as follows: since the comparison to $s$ makes $a$’s offering more attractive than $b$, $a$ is able to charge a higher price than $b$ in equilibrium. Furthermore, since $b$ has the best offering on the market, offering a highly dissimilar product from $a$ may now be detrimental to $b$ as it makes $b$’s superior value harder for consumers to see. Firm $b$ may therefore benefit from being located closer to $a$ than the dupoly equilibrium dissimilarity. These predictions yield novel policy implications. For instance, suppose that $s$ is a state-operated firm. The analysis above shows that by merely offering a product with fixed attributes, the state-operated firm may be able to engage in “standard-setting” and induce subsequent lower-cost entrants to design their products to be similar to the state-provided product, thereby reducing prices and the level of obfuscation in the market.

### v4

@@ 4.2 Tests of Preference Reversal Predictions
We run experiments mirroring the simulation exercises in Section 3.3 to test our model predictions regarding preference reversals. Specifically, we establish the presence of reversals in lottery and intertemporal choice, and test the novel model prediction that these reversals can be eliminated by manipulating the ease of comparing options to the price list.
In each domain, we construct two sets of “base" options: 3 high-risk and 3 low-risk lotteries, and 3 high-delay and 3 low-delay payoff flows (see Table 2). We elicit binary choices between these sets of options, as well as multiple price list valuations of each option. In these valuation tasks, we experimentally vary the ease of comparing each option to the price list: for lotteries, we elicit certainty equivalents as well as probability equivalents using a yardstick lottery that pays \$24 with pp% chance; for intertemporal choice, we elicit present value equivalents as well as time equivalents using a yardstick option that pays \$27.5 in $t$ days. We elicit choices and valuations for both the base options as well as a “scaled-up” version of each option, where we multiply payouts by a scale factor of 1.6.
@@
Figures 7 and 8 present results from our lottery and intertemporal choice experiments, aggregating across scale factors. Figures 7(a) and 8(a) show that binary choice rates favor the low-risk and low-delay options, with choice rates for those options well above 50%. In contrast, Figures 7(b) and 8(b) show that there is an apparent reversal in preference when the options are valued via certainty equivalents and present value equivalents: subjects assign higher valuations to high-risk and high-delay options on average. In these same figures, however, we see the predicted flipping of these valuation patterns when we manipulate the ease of comparing each option to the price list. In particular, when options are instead valued via probability equivalents and time equivalents, subjects instead assign higher valuations to the low-risk and low-delay options, thus eliminating the apparent reversal. [^26] Butler and Loomes, (2007) document a related result in lottery choice: examining inconsistency between direct choice and probability equivalents, they find that subjects are more likely to state higher PEs for the low-risk lottery yet choose the high-risk lottery in direct choice than to exhibit the opposite inconsistency. We find similar patterns in our data.
@@ 4.3 Tests of Valuation Predictions
We run a multiple price list experiment to 1) replicate patterns of inverse-S shaped probability weighting in certainty equivalents and hyperbolic discounting in present value equivalents, and 2) test the prediction, developed in Section 3.4, that these biases can be reversed by manipulating the units of valuation. Specifically, our model predicts that relative to the certainty equivalents of binary lotteries, probability equivalents of certain payments should exhibit underweighting of low probabilities and overweighting of high probabilities. Likewise, relative to the discounting revealed by the present value equivalents of delayed payments, time equivalents of immediate payments should exhibit overvaluation of delays close to the present and undervaluation of delays far in the future.
@@
Lottery Valuations. We elicit certainty equivalents $CE(l)$ for a simple lottery $l=(\overline{w},p_{l})$, and relate the normalized valuations $CE(l)/\overline{w}$ to payout probabilities $p_{l}$. We also elicit probability equivalents $PE(c)$ of a certain payment $c=(w_{c},1)$ against a yardstick lottery $(\overline{w},p)$, and relate the normalized payouts $w_{c}/\overline{w}$ to probability equivalents $PE(c)$. In our experiment, we draw $\overline{w}$ from $\{\$9,\$18,\$27\}$. For certainty equivalents, we draw $p_{l}$ from $\{0.03,0.05,0.10,0.25,0.5,0.75,0.90,0.95,0.97\}$. For probability equivalents, we draw $w_{c}$ so that $w_{c}/\overline{w}\in\{0.033,0.056,0.11,0.25,0.5,0.75,0.89,0.944,0.967\}$.
@@
Results for lotteries are presented in Figure 9. Certainty equivalent valuations follow the standard pattern: risk-averse over low probabilities and risk-seeking over high probabilities. Consistent with the model’s predictions, the pattern reverses for probability equivalents. Relative to certainty equivalents, probability equivalents are more risk-averse over low probabilities, and more risk-seeking over high probabilities. [^27] These results are consistent with Feldman and Ferraro, (2024), who find similar patterns in an expert sample of commercial agricultural producers.
@@ 4.4 Benchmarking Model Performance
Whereas the predominant approach in behavioral economics is to model non-standard elements in the DM’s value function, this paper seeks to model an orthogonal feature of choice: how difficult the DM finds a choice, given her objectives. To study whether our theory indeed explains variation in choice that is not captured by existing models, we compare the explanatory power of these modeling approaches in our binary choice data.
First, we structurally estimate our choice model, which is parameterized by a transformation $G$ that maps the value-dissimilarity ratio into choice probabilities and (in the case of lottery and intertemporal choice) domain-specific preference parameters. Across all domains, we use the same specification of $G$ developed in Section 2.4. We then compare performance against a set of models specifying the DM’s value function, which we fit to our data using logit errors. In particular, we estimate a “standard" benchmark model: a distortion-free logit model in multiattribute choice, exponential discounted utility in intertemporal choice, and expected utility in lottery choice. We also estimate a set of behavioral models: salience-weighting (Bordalo et al.,, 2013), focusing (Kőszegi and Szeidl,, 2013), and relative thinking (Bushong et al.,, 2021) in multiattribute choice; quasi-hyperbolic and hyperbolic discounting in intertemporal choice; and simplicity theory (Puri,, 2025) and cumulative prospect theory in lottery choice. Estimates are obtained using maximum likelihood; see Appendix E for details on each specification.
@@
Figure 11: Variance explained of estimated models and completeness benchmarks. $R^{2}$ values are observation-weighted. Number of free model parameters in parentheses. “LGT”, “RT”, “L1-C” refer to the Distortion-Free, Relative Thinking, and $L_{1}$-complexity models described in Appendix E.1. “EDU”, “HDU”, and “CPF-C” refer to the Exponential Discounting, Hyperbolic Discounting, and CPF-complexity models desdribed in Appendix E.2. ‘EU", “CPT", “EV CDF-C", and “EU CDF-C" refer to the Expected Utility, Cumulative Prospect Theory, risk-neutral CDF-complexity, and expected utility CDF-complexity models described in Appendix E.3. Completeness benchmarks are obtained using an ensemble of fitted models and a neural network; see Supplemental Appendix G for details.
@@
Completeness and Restrictiveness. Following Fudenberg et al., (2022), we establish “completeness benchmarks" in each of our three domains by training flexible predictive models to predict choice rates based on problem features; these benchmarks represent the predictable variation in choice rates that any model could hope to capture. We use an ensemble approach to form these benchmarks, described in detail in Supplemental Appendix G, where we combine parametric model predictions with those of a neural network trained on the data. The dashed lines in Figure 11 report the $R^{2}$ of these benchmarks, and Appendix Tables 5, 8, and 11 report the completeness index of each model following Fudenberg et al., (2022). Our model captures 70% of the predictable variation in our multiattribute choice data, and over 90% in lottery and intertemporal choice. Note that this high level of completeness in lottery and intertemporal choice suggests that there is limited scope for systematic heterogeneity across problems in how well the model captures behavior. If our model fit substantially worse on a class of problems in the dataset relative to other models, this would manifest in large performance gains of the neural network. However, this exercise cannot speak to heterogeneity across subjects; we characterize completeness only with reference to the best representative agent model.
@@
One concern is that the predictive power of our model may come at the cost of parsimony — i.e. our model is flexible enough to explain any dataset. To address this, we estimate the restrictiveness of each model: a measure of how well it fits synthetic data. Following Fudenberg et al., (2023), we simulate synthetic data in each domain according to a set of common restrictions imposed by the models we evaluate, and assess how well each model fits the synthetic data (see Appendix F for details). We obtain a restrictiveness measure $r\in[0,1]$, where $r=1$ means the model fits the synthetic data no better than a baseline model on average, and $r=0$ means the model perfectly fits the synthetic data. Appendix Tables 5, 8, and 11 report this measure in each domain. While our model is less restrictive than relative thinking in multiattribute choice ($r=0.45$ vs. 0.46), our model is more restrictive than cumulative prospect theory ($r=0.60$ vs. 0.53) and hyperbolic discounting ($r=0.59$ vs. 0.58) in lottery and intertemporal choice, respectively. These results suggest that the performance gains of our model do not come at substantial costs to model parsimony.